OpenStudy (anonymous):

Find the distance between (43, -15) & (29, -3).

OpenStudy (anonymous):

Use the wonderful distance formula! :) D = \[\sqrt{(x1-x0) ^{2} + (y1-y0) ^{2}}\] Where you subtract the x components and the y components. For instance, you gave me the points (43, - 15) and (29, -3). Plugging in, you get: \[\sqrt{(43- 29)^{2}+ (-15 - (-3))^{2}}\] = 18.439.

OpenStudy (anonymous):

its great using formula but its easy when u understand the point of it and do it making triangle on graph. as you dont have to remember the formula all the time